# IPO: Iterative Preference Optimization for Text-to-Video Generation

### v3

@@ 3.3.3 Pairwise Optimization
$L_{\text{dpo}}(\theta)=-\mathbb{E}_{(x^{w}_{0},x^{l}_{0})\sim D,t\sim\mathcal{U}(0,T),x^{w}_{t}\sim q(x^{w}_{t}|x^{w}_{0}),x^{l}_{t}\sim q(x^{l}_{t}|x^{l}_{0})}\\ \log\sigma\left(-\beta T\omega(\lambda_{t})\left(\|\epsilon^{w}-\epsilon_{\theta}(x_{t}^{w},t)\|^{2}_{2}-\|\epsilon^{w}-\epsilon_{\text{ref}}(x_{t}^{w},t)\|^{2}_{2}\right.\right.\\ -\left.\left(\|\epsilon^{l}-\epsilon_{\theta}(x_{t}^{l},t)\|^{2}_{2}-\|\epsilon^{l}-\epsilon_{\text{ref}}(x_{t}^{l},t)\|^{2}_{2}\right)\right),$ (8)
where $x^{w}$ denotes the preferred sample, $x^{l}$ the rejected sample, and $\omega(\lambda_{t})$ is a weighting function [11, 14, 27].
During training, we use a selective sampling strategy to construct high-quality preference pairs by selecting the highest-ranked (preferred) and lowest-ranked (rejected) samples based on critic model evaluations. This ensures clearer preference distinctions and enhances learning. These pairs $(y^{w},y^{l})$ form the final training data.
The optimization objective combines the Diffusion-DPO loss with two auxiliary negative log-likelihood (NLL) terms. The first, weighted by $\lambda_{1}{=}0.2$, regularizes the preferred samples to align with human preferences. The second, weighted by $\lambda_{2}{=}0.1$, anchors the model to real video distributions to prevent overfitting. The final loss is:
$\mathcal{L}(\theta)$ $=L_{\text{dpo}}(\theta)$ (9)
@@
$\quad+\lambda_{1}\cdot\mathbb{E}_{(\textbf{{y}},x)\sim\mathcal{D}^{\mathrm{sample}}}\bigl{[}-\log p_{\theta}(y^{w}|x)\bigr{]}$
@@
$\quad+\lambda_{2}\cdot\mathbb{E}_{(\textbf{{y}},x)\sim\mathcal{D}^{\mathrm{real}}}\bigl{[}-\log p_{\theta}(y|x)\bigr{]},$
where $\mathcal{D}^{\mathrm{sample}}$ represents the set of preference pairs derived from the critic model’s evaluation, and $\mathcal{D}^{\mathrm{real}}$ denotes the real video data used for regularization.
3.3.4 Iterative Optimization
Iterative optimization is advantageous for complex tasks like preference alignment in reinforcement learning. Rather than approximating the optimal policy $\pi_{\theta}$ in a single step, which is computationally challenging, it decomposes the problem into manageable stages. This stepwise refinement enables incremental improvements, progressively aligning the model with human preferences in a more efficient and controlled manner. A key benefit of iterative optimization is that it enables more stable training. Rather than attempting to directly approximate the reference distribution $\pi_{\text{ref}}\exp\left(\frac{1}{\beta}r(x_{0},c)\right)$ in a single step, the process proceeds in multiple stages. In each iteration, we approximate the distribution using the following formula:
$\pi_{\text{ref}}^{(k)}=\pi_{\text{ref}}^{(k-1)}\exp\Big{(}\frac{r^{k}}{\beta_{k}}\Big{)},$ (10)
where, $r^{k}$ denotes the reward function at the $k$-th iteration, and $\beta_{k}$ is a hyperparameter controlling the optimization rate at each stage. We initialize $r^{0}=0$ to start with a neutral reward function. As iterations progress, the reward function is refined, guiding the model toward an optimal policy.
Iterative refinement improves training control, enabling the model to gradually align with human feedback while avoiding early overfitting or divergence. By utilizing intermediate distributions at each stage, the method accommodates more complex feedback signals, leading to a more robust model that generalizes well across diverse real-world scenarios. Additionally, iterative optimization mitigates issues like catastrophic forgetting and instability, which can occur when attempting to approximate the optimal policy in a single step. Each iteration builds on the previous one, ensuring stable and reliable convergence toward the target policy. In this way, iterative preference optimization (IPO) enhances preference-based optimization by providing better training control, faster convergence, and improved generalization, making it a valuable tool for optimizing policies based on human feedback.
Method Quality Semantic Total
Score Score Score
CogVideoX-2B 82.18 75.83 80.91
CogVideoX-5B 82.75 77.04 81.61
CogVideoX-2B-IPO-3 83.92 78.00 82.74
Table 1: Comparison of IPO with baselines CogVideoX models.
Models Subject Motion Aesthetic
Consist. Smooth Quality
CogVideoX-2B 96.78 97.73 60.82
w/o real data sft 96.56 97.81 61.25
w/ real data sft 96.77 97.84 61.55
Table 2: Ablation studies on effects of real data for finetuning IPO.
@@ 4 Experiments
Dynamic
Degree
Image
Quality
CogVideoX-2B 80.91 82.18 75.83 96.78 96.63 98.89 97.73 60.82 59.86 61.68
CogVideoX-2B-IPO1 81.69 82.87 77.01 96.77 97.01 99.01 97.84 61.55 64.33 62
CogVideoX-2B-IPO2 82.42 83.53 77.97 96.81 97.23 99.21 98.03 62.35 68.76 61.77
CogVideoX-2B-IPO3 82.74 83.92 78.00 96.79 97.48 99.35 98.17 62.31 69.46 62.87
Table 3: Ablation studies on the iteration number adopted in IPO. The evaluation is performed on VBench.
Figure 4: Qualitative Comparison of our proposed IPO model and the baseline CogVideoX-2B model. We can see IPO improves the baseline in the aspects of prompt following (the 1st line), aesthetic quality (the 2nd line), subject consistency (the 3rd line) and motion smoothness (the 4th line). This demonstrates the efffectiveness of our IPO for enhancing the generation capability of video found models. Additional visualization results are available in the supplementary materials.
4.1 Experimental Setup
4.1.1 Critic Model Training
To unify paired data sorting and individual video scoring, we fine-tune the pre-trained VILA [19], leveraging its strengths in video comprehension and instruction following. Key modifications include integrating reasoning chains before scoring or ranking and enhancing decision-making with explicit thought processes. This improves result accuracy and the quality of natural language feedback.
[…]
4.1.2 T2V Alignment Training
[…]
4.2 Results
4.2.1 Quantitative Results
Quantitative results are evaluated using the VBench metric, which measures video generation quality in terms of temporal coherence, semantic alignment, and overall plausibility. As shown in Tab. 1, our experiments reveal that the optimized $2B$ model outperforms the baseline $5B$ model across all metrics. The optimized model shows significant improvements in temporal coherence, enhancing frame transitions and minimizing abrupt changes. It also demonstrates stronger semantic alignment, better capturing the intended meaning and context of input prompts. These results highlight the effectiveness of our post-training strategy in addressing key limitations of the baseline model. The overall VBench score further confirms that the optimized $2B$ model achieves superior video quality while maintaining computational efficiency, making it a practical solution for optimizing smaller-scale models.
4.2.2 Qualitative Results
To further assess the benefits of our approach, we provide qualitative comparisons between the baseline and optimized models. Visual results, shown in Fig. 4, reveal that the optimized $2B$ model generates videos with better semantic fidelity, accurately capturing the intent of input prompts and their context. It also ensures improved temporal consistency by minimizing artifacts and sudden transitions between frames, common in the baseline outputs. Additionally, the optimized videos exhibit enhanced realism, with fewer incoherent structures and greater plausibility, making them more visually appealing and aligned with human perception. This further demonstrate the effectiveness of IPO.
4.3 Ablation Studies
4.3.1 Incorporating Real Video Data
To explore the impact of real video data during post-training, we integrated real samples into the training process and applied the original diffusion model loss with reinforcement learning objectives. This modification greatly improves generation quality by enhancing temporal coherence, reducing flickering, and strengthening frame consistency. Furthermore, the inclusion of real-world examples boosts semantic alignment, guiding the model to generate more realistic motion dynamics and contextual details.
As shown in Tab. 2, the introduction of real video data (indicated by the inclusion of sft loss) leads to measurable improvements across multiple VBench metrics, including temporal flicker, motion smoothness, and aesthetic quality. Notably, the optimized model achieves higher scores in semantic consistency and dynamic degree, indicating a significant enhancement in the realism and coherence of generated videos. These results underscore the effectiveness of integrating real-world data with reinforcement learning to refine video generation outputs, enabling the model to produce videos with superior visual and temporal fidelity.
Method Quality Semantic Total
Score Score Score
@@ 4.3.1 Incorporating Real Video Data
CogVideoX-2B-KTO1 82.47 83.56 78.08
CogVideoX-2B-DPO1 82.42 83.53 77.97
Table 4: Ablation studies on impacts of different preference optimization strategies for IPO: DPO and KTO.
4.3.2 Comparison of different RL Methods
To evaluate the performance of the proposed online reinforcement learning framework, we conducted experiments using two different feedback alignment methods: Knowledge Transfer Optimization (KTO) and Direct Policy Optimization (DPO). The results, presented in Tab. 4, demonstrate that our framework consistently improves video generation quality across multiple metrics, regardless of the specific reinforcement learning method employed. All evaluation metrics across different dimensions are provided in the supplementary materials, further supporting this comprehensive improvement.
Both KTO and DPO lead to significant enhancements over the baseline CogVideoX-2B model, showing improvements in temporal flicker reduction, motion smoothness, and dynamic degree. These results validate the generality and robustness of the proposed online reinforcement learning framework, which effectively integrates human feedback to optimize video generation outputs. For example, the inclusion of either KTO or DPO raises the Aesthetic Quality and Dynamic Degree metrics compared to the baseline model, underscoring the benefits of aligning the generated outputs with human preferences.
Between the two methods, KTO achieves consistently better performance than DPO across most metrics. Notably, KTO yields a higher Dynamic Degree score (68.83 vs. 68.76) and improved Motion Smoothness (97.93 vs. 98.03), suggesting that KTO is more effective at producing realistic and temporally coherent videos. Additionally, KTO demonstrates greater stability during training, as evidenced by its higher overall scores across Subject Consistency and Background Consistency, which measure adherence to semantic and contextual details. These findings highlight the advantages of the KTO method in leveraging multi-step optimization, enabling the generation of more plausible and visually consistent video outputs compared to DPO. This confirms our IPO is flexible to adapt to different optimization methods. In addition, KTO exhibits superior performance and stability, making it a preferred choice.
4.3.3 Effect of Multi-Round Training
The impact of multi-round reinforcement learning is assessed by training models with 1, 2, and 3 rounds of iterative optimization, where a single round represents the traditional offline post-training approach. Results in Tab. 3 show that performance improves with each additional round, as multi-round training better aligns the model with human feedback and refines generation quality. These findings confirm that iterative reinforcement learning effectively enhances both temporal and semantic consistency.
Figure 5: Human evaluation of IPO and CogVideoX-2B.
4.4 Human Evaluation
Fig. 5 presents the results of a human evaluation comparing the IPO model with the baseline CogVideo-2B The result shows that our IPO outperforms the baseline in all dimensions. In the ”Consistency” dimension, the two models performed equally in around 40.5 percent of the cases, indicating that the baseline excels in maintaining text consistency. However, in the ”Motion” and ”Faithfulness” dimensions, the baseline’s performance is average, while the IPO model demonstrates a significant improvement, particularly in generating more natural motion and better alignment with the given text. All evaluation metrics across different dimensions are provided in the supp. materials

### v4

@@ 3.3 Iterative Alignment for Video Generation
$L_{\text{dpo}}(\theta)$ $=-\mathbb{E}_{(x^{w}_{0},x^{l}_{0})\sim D,\,t\sim\mathcal{U}(0,T)}\log\sigma\!\Big(-\beta T\omega(\lambda_{t})\cdot\big[$ (6)
$\|\epsilon^{w}-\epsilon_{\theta}(x_{t}^{w},t)\|^{2}-\|\epsilon^{w}-\epsilon_{\text{ref}}(x_{t}^{w},t)\|^{2}-(\|\epsilon^{l}-\epsilon_{\theta}(x_{t}^{l},t)\|^{2}-\|\epsilon^{l}-\epsilon_{\text{ref}}(x_{t}^{l},t)\|^{2})\big]\Big).$
where $\omega(\lambda_{t})$ is a time-dependent weighting function, and $\epsilon^{w}$, $\epsilon^{l}$ are noise vectors from the forward process. To improve generalization, we add two auxiliary negative log-likelihood (NLL) terms:
@@
$\mathcal{L}_{\text{total}}^{\text{DPO}}(\theta)$ $=L_{\text{dpo}}(\theta)+\lambda_{1}\cdot\mathbb{E}_{\mathcal{D}^{\mathrm{sample}}}\bigl[-\log p_{\theta}(y^{w}|x)\bigr]+\lambda_{2}\cdot\mathbb{E}_{\mathcal{D}^{\mathrm{real}}}\bigl[-\log p_{\theta}(y|x)\bigr],$ (7)
where $\mathcal{D}^{\mathrm{sample}}$ contains preference samples, and $\mathcal{D}^{\mathrm{real}}$ contains real videos for regularization.
For pointwise alignment, we incorporate Diffusion-KTO Li et al. (2025), which optimizes a nonlinear utility-based objective without requiring preference pairs. The core loss is defined as:
$L_{\text{kto}}(\theta)=\mathbb{E}_{x_{0}\sim\mathcal{D},\,t\sim\text{Uniform}(0,T)}\left[U\left(w(x_{0})\left(\beta\log\frac{\pi_{\theta}(x_{t-1}|x_{t})}{\pi_{\text{ref}}(x_{t-1}|x_{t})}-Q_{\text{ref}}\right)\right)\right],$ (8)
where $w(x_{0})\in\{\pm 1\}$ denotes desirability, and the reference value $Q_{\text{ref}}$ is:
$Q_{\text{ref}}=\beta\cdot\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(a|s)\,||\,\pi_{\text{ref}}(a|s)\right].$ (9)
To stabilize training, we further regularize the loss using real-world data as:
@@
$\mathcal{L}_{\text{total}}^{\text{KTO}}(\theta)=L_{\text{kto}}(\theta)+\lambda\cdot\mathbb{E}_{\mathcal{D}^{\mathrm{real}}}\bigl[-\log p_{\theta}(y|x)\bigr].$ (10)
Note that these two optimization paths are activated independently within our optimization proces according to the available preference data format, allowing flexible deployment under varied alignment strategies. Our dual-iterative optimization framework offers several key advantages: 1) we dynamically align the reward model drift and generator distributional shifts through dual optimization, improving training robustness and stability; 2) we significantly reduce requirements on large-scale human annotations by leveraging critic self-refinement; 3) we provide a unified yet flexible framework for handling both pointwise and pairwise preference data; and 4) we achieve stronger generalization and stability by continuously aligning generation behavior with up-to-date human preferences.
@@ 4 Experiments
4.1 Implementation Details
We fine-tune pre-trained VILA Lin et al. (2023) (13B/40B) models as reward models, leveraging their strong video-text understanding and instruction-following abilities. A small set of Chain-of-Thought (CoT) annotations is used for warm-start, followed by 5 epochs of training (LR $2\text{e-}5$, batch size 4 on 32 GPUs). Videos are represented by 16 uniformly sampled frames with padding-based preprocessing. Other settings follow VILA defaults. We further apply Self-Refined Preference Optimization (SRPO) with CoT-guided inference, self-consistency voting, and confidence filtering, generating 20K pseudo-labels per iteration for iterative fine-tuning.
Figure 2: Qualitative Comparison of our proposed Dual-IPO model and the baseline CogVideoX-2B model. We can see Dual-IPO improves the baseline in the aspects of prompt following (the 1st line), aesthetic quality (the 2nd line), subject consistency (the 3rd line) and motion smoothness (the 4th line). This demonstrates the efffectiveness of our Dual-IPO for enhancing the generation capability of video found models. Additional visualization results are available in the supplementary materials.
T2V alignment training.
For generator alignment, the critic assesses semantic fidelity and temporal consistency under the Dual-IPO framework with Diffusion-DPO and Diffusion-KTO. Training uses a batch size of 512 and LR $2\text{e-}6$ (raised to $2\text{e-}5$ after stabilization) across 128 GPUs. For each caption, the T2V model generates 8 candidates, with the top-1 and bottom-1 (by critic ranking) forming preference pairs. Each iteration constructs about 100K such pairs, and a single training round takes on average two weeks.
Datasets.
To construct large-scale and diverse training data, we first build a pool of structured elements (subjects, attributes, spatial relations, actions), which are randomly composed and expanded into natural language captions using an LLM, producing about one million captions. The target T2V model then generates corresponding videos for iterative optimization. In addition, a subset of real video-text pairs from VidGen-1M Tan et al. (2024a) is incorporated for joint fine-tuning to enhance diversity and grounding in natural distributions.
Evaluation.
For evaluation, we use both automatic metrics and human assessments. We report the overall and per-dimension scores from the VBench benchmark Huang et al. (2024), including semantic consistency, motion smoothness, and visual quality. To complement these metrics, we conduct user studies to verify synthesis quality.
4.2 Experimental Results
A customized reward model is necessary for reliable alignment.
Table 1: Comparison of reward models by human preference accuracy and downstream VBench score after DPO training.
Reward Models Accuracy (%) VBench
VideoScore 63.58 80.87
VideoAlign 65.21 81.27
VideoReward 68.44 81.31
Ours 81.33 81.54
We first evaluate the effectiveness and generalizability of our reward model by measuring alignment with human preference labels. We randomly sample 1,000 videos from CogVideoX-2B and WanX-1.3B, and manually annotate them as a reference set. Our critic is compared against three open-source models: VideoScore He et al. (2024), VideoAlign Liu et al. (2025a), and VisionReward Xu et al. (2024). As shown in Tab. 1, our reward model achieves 81.33% accuracy, significantly outperforming VisionReward (68.44%), VideoAlign (65.21%), and VideoScore (63.58%). The weaker generalization of these models is due to their training on different data and annotation schemes.
We further use each critic to construct preference datasets for fine-tuning CogVideoX-2B via Diffusion-DPO. As shown in Tab. 1, our critic boosts VBench from 80.91 to 81.54 (+0.67), while VideoScore slightly degrades performance, and VideoAlign and VisionReward yield only minor gains around +0.3. This shows critic models must be tailored to the specific T2V generator they supervise. Notably, our reward model is trained on just 6,000 annotated pairs, fewer than other models, yet achieves strong alignment. We attribute this to Chain-of-Thought reasoning and VLM-based inference, which offer richer supervision from limited data.
Comparison with current sota.
We evaluate our method on VBench (Tab.5), where our optimized 2B model surpasses the baseline 5B across all dimensions, showing that Dual-IPO post-training effectively boosts generation quality in smaller models. As shown in Tab.2, Dual-IPO consistently improves motion, aesthetics, and consistency during iterative training, demonstrating both stability and effectiveness. These gains also reflect better prompt alignment and fewer motion artifacts in generated videos.
To further validate our generalization, we apply our model to both CogVideoX-5B and Wan-1.3B. As reported in Tab. 5, after three iterations, CogVideoX-5B improves from 81.61 to 84.63 on VBench, and Wan-1.3B increases from 84.26 to 86.28. These gains are evident in both visual quality and prompt consistency. We also evaluate the final models after five Dual-IPO in Tab. 5, it turns out that our model yields consistent performance gains, demonstrating both the effectiveness and generalizability of our proposed framework across architectures and scales.
Table 2: Ablation studies on the iteration number adopted in Dual-IPO. The evaluation is performed on VBench.
@@ 4.2 Experimental Results
CogVideoX-2B 80.91 82.18 75.83 96.78 96.63 98.89 97.73 60.82
CogVideoX-2B-IPO1 81.69 82.87 77.01 96.77 97.01 99.01 97.84 61.55
CogVideoX-2B-IPO2 82.42 83.53 77.97 96.81 97.23 99.21 98.03 62.35
CogVideoX-2B-IPO3 82.74 83.92 78.00 96.79 97.48 99.35 98.17 62.31
Table 3: Dual-IPO under various baselines.
Model Total Quality Semantic
@@
CogVideoX-2B-IPO-3 82.74 83.92 78.00
CogVideoX-5B 81.61 82.75 77.04
CogVideoX-5B-IPO-3 84.63 85.40 81.54
Wan-1.3B 84.26 85.30 80.09
Wan-1.3B-IPO-3 86.28 86.38 85.87
Table 4: Ablation studies: DPO vs KTO.
Method Total Quality Semantic
CogVideoX-2B 80.91 82.18 75.83
CogVideoX-2B-KTO3 82.47 83.56 78.08
CogVideoX-2B-DPO3 82.42 83.53 77.97
Table 5: Comparison with sota T2V models on VBench.
Model Total Quality Semantic
Vidu Q1 87.41 87.28 87.94
Open-Sora-2.0 84.34 85.40 80.12
Sora 84.28 85.51 79.35
CausVid 84.27 85.65 78.75
Wan2.1 86.22 86.67 84.44
CausVid 83.88 85.21 78.57
HunyuanVideo 83.24 85.09 75.82
CogVideoX-5B 82.01 82.72 79.17
+Ours 86.57 87.00 84.84
CogVideoX-2B 80.91 82.18 75.83
+Ours 82.74 83.92 78.00
Wan-1.3B 84.26 85.30 80.09
+Ours 88.32 87.83 90.25
Qualitative results.
To further assess our approach, we present qualitative comparisons in Fig. 2. The optimized model consistently produces videos with better semantic fidelity, accurately reflecting prompt intent and context. It also improves temporal consistency by reducing artifacts and abrupt transitions. Additionally, the outputs show enhanced realism with fewer structural errors, demonstrating the effectiveness of Dual-IPO.
Figure 3: Comparison of IPO and baselines: (a) VBench shows DPO degrades, while Dual-IPO improves steadily. (b) Human evaluation compares Dual-IPO, CogVideoX-2B, and Ties on consistency, motion, and faithfulness.
Table 6: Ablation study on SRPO during the third to fourth Dual-IPO iteration on the CogVideoX-2B baseline.
Setting Critic Update SRPO loss PCE VBench Score
Baseline 82.74
No update 82.33
SRPO (full) 82.91
SRPO loss 82.83
PCE 82.69
Effectiveness of Self-Refined Preference Optimization (SRPO).
To evaluate SRPO’s role in improving critic generalization, we conduct controlled ablations during the training of CogVideoX-2B in Tab. 6. Specifically, we examine model performance between the third and fourth IPO iterations under two conditions: (1) keeping the critic model fixed, and (2) updating the critic via SRPO. We can observe that without critic update, performance declines from 82.74 to 82.33, indicating overfitting or misalignment. In contrast, applying SRPO increase the performance from 82.74 to 82.91, demonstrating the necessity and efficacy of our proposed reward model in-the-loop optimization.
We ablate SRPO components. Removing PCE with SRPO loss causes minimal change (82.69), suggesting limited impact. Replacing SRPO loss with DPO, even with PCE, drops performance to 82.83, confirming SRPO loss is crucial for critic robustness and signal quality. These results underscore the importance of reward modeling, critic refinement, and self-refinement for stable alignment.
4.3 Effectiveness of Iterative Alignment
Why multi-round iteration matters.
To evaluate iterative optimization, we compare Dual-IPO with a static baseline using Diffusion-DPO without updating the preference data or the reward model. As shown in Fig.3a, the static baseline improves briefly but degrades due to overfitting and misaligned rewards, while Dual-IPO maintains steady performance gains with progressive data improvement and reward optimization. Tab.5 and Tab. 2 confirm that Dual-IPO outperforms the static method in overall VBench scores and all quality dimensions, with consistent improvements in motion, aesthetics, and semantic consistency.
Benefits of incorporating real video data.
Table 7: Ablation study on the effects of incorporating real video data during Dual-IPO fine-tuning.
Model Total Quality Semantic
CogVideoX-2B 80.91 82.18 75.83
only real data sft 80.63 81.79 76.02
w/o real data sft 81.17 82.33 76.53
w/ real data sft 81.69 82.87 77.01
We examine whether real-world video data improves post-training. Using VidGen-1M Tan et al. (2024a) with diffusion loss and preference optimization, we observe consistent gains in total, quality, and semantic scores (Tab. 7). In contrast, real data alone slightly hurts performance, suggesting it is most effective when combined with preference guidance. We leave further analysis to future work.
Comparing DPO and KTO within Dual-IPO.
To assess flexibility, we compare DPO and KTO within Dual-IPO (Tab. 5). Both outperform the baseline. KTO slightly improves Total and Semantic scores, while DPO leads in Quality, showing Dual-IPO supports diverse alignment. KTO’s more stable semantic gains suggest better high-level consistency, confirming the adaptability of Dual-IPO and effectiveness of KTO for fine-grained optimization.
Human evaluation.
Fig. 3b shows the results of human visual evaluation that compare our Dual-IPO with the CogVideo-2B baseline. Obvisouly, Dual-IPO outperforms the baseline across all dimensions. In “Consistency,” both models perform equally in  40.5% of cases, suggesting the baseline handles text alignment reasonably well. However, Dual-IPO shows clear gains in “Motion” and “Faithfulness,” generating more natural movement and better text alignment. Full evaluation metrics are in the appendix.